# The impact of medium-width bands on the selection, and subsequent luminosity function measurements, of high-z galaxies

### v1

@@ Abstract
New, ultra-deep medium-width photometric coverage with JWST’s NIRCam instrument provides the potential for much improved photo-z reliability at high redshifts. In this study, we conduct a systematic analysis of the JADES Origins Field, which contains 14 broad- and medium-width near-infrared bands, to assess the benefits of medium band photometry on high-z sample completeness and contamination rates. Using imaging with depths of AB mag , we conduct an experiment to observe how high-z selections differ when images are artificially degraded or bands are removed. In parallel, the same experiments are conducted on simulated catalogues from the JAGUAR semi-analytic model to examine if the behaviour from observations can be replicated. We find sample completeness is high () and contamination low () when in the regime, even without the use of any medium-width bands. The addition of medium-width bands leads to notable increases in completeness () but multiple bands are required to improve contamination rates due to the small redshift ranges over which they probe strong emission lines. Incidents of Balmer-Lyman degeneracy increase in the regime and this can be replicated in both simulated catalogues and degraded real data. We measure the faint-end of the UV LF at , finding high number densities that agree with previous JWST observations. Overall, medium bands are effective at increasing completeness and reducing contamination, but investment in achieving at least comparable depths in the blue (m) as achieved in the red is also found to be key to fully reducing contamination from high-z samples.
@@
In the first few years since the successful commissioning of the James Webb Space Telescope (JWST), the study of the first billion years of the Universe’s history has undergone a revolution. Exquisitely deep, high resolution imaging and spectroscopy has resulted in the photometric selection of ’s of high-z galaxies ( e.g. Harikane et al., 2023; Adams et al., 2024; Finkelstein et al., 2024; McLeod et al., 2024; Donnan et al., 2024, D’Silva et al. In Prep) and spectroscopic confirmation of ’s (e.g. Bunker et al., 2024; D’Eugenio et al., 2024; Price et al., 2024; Mascia et al., 2024; Heintz et al., 2024). These observations have revealed higher number densities of UV luminous galaxies than first expected from extrapolated HST observations (Oesch et al., 2018; Harikane et al., 2022), pushed the redshift frontier to (Curtis-Lake et al., 2023; Wang et al., 2023; Carniani et al., 2024) and discovered previously unknown classifications of sources (e.g. little red dots Labbé et al., 2023; Matthee et al., 2023; Kokorev et al., 2024).
However, even though JWST provides substantially more information on the Universe than we previously had, there have been cases where robust looking sources at very high redshifts have turned out to be sample contaminants at lower redshift (Arrabal Haro et al., 2023). Degeneracies between the Lyman and Balmer breaks in a galaxy’s spectra and the chance alignment of strong emission lines, in particular, can result in cases where low-z galaxies enter photometric samples of high-z galaxies (See also discussion in Zavala et al., 2023; Gandolfi et al., 2025, for dusty sources). While spectroscopy is the most powerful tool to obtain precise and accurate redshift measurements, JWST has begun to produce deep fields exceeding AB magnitudes of 30 in depth. In this faint regime, follow-up spectroscopy becomes increasingly time consuming, bordering infeasible in fact due to the anticipated costs to obtain required signal-to-noise ratios (Giardino et al., 2022). The use of medium-width photometric bands, while not producing the level of detail as spectroscopy, may provide a route to increase the confidence of high redshift solutions for ultra-faint sources with smaller overall costs. In Cycle 2, the accepted General Observer programme PID 3215 (PI: D. Eisenstein Eisenstein et al., 2023b) produced ultra-deep imaging in multiple medium with bands over a region the GOODS-South field now dubbed the JADES Origins Field (JOF), reaching AB magnitudes of 30.3 in broad- and medium-width bands (0.32as diameter apertures), providing a total of 14 NIRCam bands compared to the more standard 7-8 that many other Cycle 1 and 2 programmes utilised.
@@ 3 Galaxy Selection Procedures
A catalogue of the sources in the Jades Origins Field is produced using the SExtractor software in dual image mode. Here, a weighted stack of the red broad band images (F277W, F356W, F444W) is used at the selection band. Aperture photometry is measured in both 0.2 and 0.32 arcsecond apertures. A set of science images PSF homogenised to the F444W band are also produced using empirical PSF models generated following the technique used in Weaver et al. (2024). We find the profile of the PSF model is similar in shape but slightly broader than theoretical models generated by WebbPSF (Perrin et al., 2012, 2014). In the following subsections we describe how we select and measure the galaxies using different criteria and later how the results depend on these.
@@ 3.1 Photometry and Photo-Z procedure
Photometric redshifts for our sources are conducted with the use of the EAZY-py SED fitting tool (Brammer et al., 2008). Here, we employ the use of the default FSPS templates as well as Sets 1 and 4 generated in the work by (Larson et al., 2023). These additional templates provide galaxy spectra which are younger, bluer and have stronger emission lines compared to the default template set on its own. EAZY is run using the Madau (1995) treatment for absorption by neutral Hydrogen, adds dust attenuation up to 3.5 magnitudes from Calzetti et al. (2000) and employs no galaxy luminosity prior. EAZY is run twice with upper redshift limits of 25 and 6 in order to conduct a direct comparison between the best fitting high-z and low-z solutions. We also investigate how our results would change when using different size apertures.
@@ 3.3 The initial full sample of high-z candidates
Utilising the full set of NIRCam data available in this field and conducting the above photo-z selection procedure, we obtain 50 galaxies which we class as robust at using the full JOF dataset. The distribution of sources is as follows, with a full table of high-z candidates presented in Table LABEL:tab:fulllist: 27 sources with , 12 sources with , 7 sources with , 2 sources with and 1 source with . Redshifts and values for the sample are shown in Figure 4. Since we only utilise JWST data in our selections, our sample is incomplete at due to the demand for a complete break to be observed and the lack of data bluewards of 0.9 microns. While HST data is present in this field, it is significantly shallower (mag 28.3-29.1 in F606W) and non-uniform (Whitaker et al., 2019). Since ultra-deep HST data can not always be a guarantee for such JWST fields, we focus on the use of just JWST photometry for the purpose of this study. Below we discuss some of the details of our selected galaxies and comparisons to previous similar studies on this field.
@@ 3.3.1 A robust candidate galaxy at
The highest redshift, robust galaxy in our catalogues that passes visual inspection is a faint () galaxy which is selected at (See Figure 5) with a of 15 relative to a low redshift solution at . This agrees with the expectations set out in the JOF survey paper (Eisenstein et al., 2023b). Key for determining this is the slightly blue slope measured in the detected bands and the lack of line emission in the F250M band, which covers H and O[III] at . This object is also identified in Robertson et al. (2024) as their highest redshift robust candidate with , meaning this source has now been selected as robust with very similar photometric redshifts in two completely independent works. NIRSpec spectroscopy was taken of this source, but due its faint luminosity and the large slit losses in the observations, no definitive redshift was obtained (Carniani et al., 2024).
@@
JADES+53.09731-27.84714 | Robust | 11.53 | Faint, Broad PDF | ||
@@
JADES+53.04017-27.87603 | Robust | 12.10 | Faint | ||
JADES+53.03547-27.90037 | Robust | 12.38 | Faint | ||
JADES+53.06475-27.89024 | Robust | 12.93 | Robust | ||
@@
JADES+53.08294-27.85563 | Robust | 14.39 | Balmer Break, Blended, (Spec-z 14.32) | ||
@@
JADES+53.07248-27.85535 | Contributors | 10.61 | Faint (0.32), Robust (0.2) | ||
JADES+53.03139-27.87219 | Contributors | 10.76 | … | Faint | |
JADES+53.09292-27.84607 | Contributors | 11.05 | Faint, low-z 1.6 micron bump, Masked | ||
@@
JADES+53.07385-27.86072 | Secondary | 13.06 | Balmer Break |
@@ 3.4 Comparisons to the sample selected in Robertson+24
For a full breakdown of our findings for the sources presented in Robertson et al. (2024), we find that: 4/9 sources in their main sample are selected in our work, 0/4 of their lower redshift sources are selected and 0/2 of their dubious sources are selected. In total, four sources are identified in our work to be at and better fit by Balmer breaks than Lyman-breaks. These include both of the sources described as dubious in the work by Robertson et al. (2024). The additional two include one of their lower-z candidates and the, now spectroscopically confirmed, source which is blended with a neighbour resulting in detections in the bluest bands. We find one source is better fit at as opposed to , however this source has a redshift PDF with a large positive wing and so it is not in significant large tension with the initial Robertson et al. (2024) estimations.
@@
Various studies targeting high redshift samples have primarily made use of two different sized circular apertures when conducting initial flux measurements. These have been 0.3-0.32 arcsecond diameters (amounting to 5-6 pixels in radius) or 0.2 arcsecond diameters (amounting to 3.3 pixels in radius). There are potential advantages and disadvantages to using each. The use of larger apertures is slightly more forgiving for the precision of the WCS alignment between photometric bands and it encloses more flux from the PSF (), requiring smaller aperture corrections. The smaller aperture size can enable higher SNR measurements to be made but can be more greatly influenced by bad pixels, correlated noise, minor image misalignments and it requires larger corrections for flux outside of the apertures. At a radius of 0.1as, the PSF for JWST’s reddest bands has a very steep slope for its encircled flux verses radius, meaning precise PSF models are required as small changes to the FWHM will result in larger changes in the aperture correction verses the use of a larger aperture.
To assess if there are significant differences in sample selection as a consequence of this choice, we repeat our selection procedure on a catalogue using 0.2 arcsecond apertures. This catalogue selects 76 sources at compared to the 50 sources that the 0.32as catalogue did. Of these sources, 46 are consistent between the two samples. The primary reason sources are lost are due to the poorer of the fit, with only one source gaining a significant detection in blue bands when smaller apertures are used. This source was at and in the regime where pure JWST selections become incomplete due to the positioning of the Lyman break within the F090W band. Of the 30 sources gained, 50% (15/30) were below the regime of the 0.32 arcsecond catalogue and 30% (9/30) were in the 5-7 regime in the previous 0.32 arcsecond catalogue. Under the assumption these redshifts are correct, this amounts to a completeness in the 5-7 regime in this field when using 0.32as apertures. This completeness is lower than anticipated from our simulation work which we conduct in Section 4.2, however it is dependent on a very small number of sources. We conduct more comparisons between these catalogues in later sections, such as seeing if they produce a consistent UV Luminosity Function. Additional objects selected with the 0.2as aperture catalogue are also included in Appendix A alongside our fiducial catalogue.
@@
Among the faint sources added to the high-z sample are an additional two galaxies from the Robertson et al. (2024) sample. Non-selected galaxies from that work with high-z solutions typically only just fail our detection criteria (e.g. 4.5 in rest-frame UV). The spectroscopically confirmed galaxy at remains non-selected even with the smaller aperture size on account of contamination from the neighbouring source producing a 2.4 flux in F115W.
The use of smaller apertures also reveals three new ultra-high redshift sources at in our selections. One at is immediately rejected by eye as likely being a shredded segment of a larger resolved source. The source labelled as ID 7891 () in Appendix A has dubious F300M and F444W fluxes which draws suspicion over whether the source is a Balmer break at with H/O[III] and H impacting band measurements around 2.5 and 3.5 microns respectively, as F300M and F444W indicate a lower potential continuum level than fit. The source labelled as ID 908 () has detections in 4/7 of the bands it is detected in and detections in 3/7 bands, making its precise SED shape relatively uncertain, particularly redwards of 4 microns. Consequently, the source presented in Figure 5 remains the most robust looking ultra-high-z source in the sample and it is selected with both choices of aperture size. With the smaller aperture size, the to a low-z solution increases from 15 to 35 for this source.
@@ 4.1 Adding and Removing Bands from Real Data
Our original selection using the full dataset (C-ALL) identified 50 galaxies at . When applying our selections to the three catalogues with fewer medium bands (C-0, C-1, C-2), each respectively identified 46, 51, 51 candidates. Of these objects, 41, 44, 45 objects are consistent with the fiducial sample (C-ALL), showing that sample overlap with C-ALL is relatively high (82-90%) and increases with increasing numbers of photometric bands. The largest increase in sample overlap occurs when a single medium-band is added to the broad-band-only dataset. Sources which are identified as confident in C-ALL but omitted when fewer bands are used are consistent between the different band combinations (e.g. sources selected in C-ALL but omitted in C-2 were also omitted in C-1 and C-0).
We begin our analysis by examining the 9 objects omitted from selection when comparing the C-ALL catalogue to the C-0 catalogue. Half (5) of these sources are low SNR, with F200W detections at . The remaining 4 are between 8 and 12. All objects retain initial high-z solutions in C-0, C-1 and C-2 but they fail requirements regarding significant secondary low-z peaks, indicating these sources are lost from the sample due to inadequate confidence regarding the Lyman and Balmer break degeneracy. This degeneracy is reduced when medium-width bands reveal a lack of optical emission lines. One of the sources lost is JADES+53.02868-27.89301 from Robertson et al. (2024), here we find this object is only selected in the C-2 or C-ALL catalogues and not in C-0 or C-1. This both highlights the potential benefits that multiple medium bands can provide high-z selection and also shows that we should not entirely overlook lower confidence sources in other surveys which may be unselected as a consequence of expected incompleteness levels (In this case, a - candidate could have been overlooked if located in another survey programme).
@@
We next examine sources that are added to the high-z sample when fewer bands are used. For the C-0, C-1 and C-2 catalogues we find that 5, 7, 6 new sources are selected compared to the C-ALL sample. These extra sources are largely consistent between these catalogues, with 3/5 and 6/6 sources in the C-0 and C-2 catalogues appearing in the C-1 catalogue. This means there are 9 unique sources which get selected as high-z when fewer photometric bands are available. We find that the sources that are selected outside of the C-ALL catalogue are faint, with 6/9 sources below in the rest frame UV. When examining these new sources in the C-ALL catalogue, we find 2/9 sources have primary photo-z solutions that are low-z Balmer breaks whilst all others have similar high-z primary solutions but do not meet cuts relative to low-z solutions. Overall, the inclusion of many medium-width bands reveals that these faint sources are more uncertain than first thought with the sole use of wide bands.
There are two out of the nine sources that are not selected in the C-ALL catalogue which are potentially legitimate high-z sources. One source (the brightest one) is a legitimate looking candidate which scatters above our selection only in C-0. The second source may be potentially legitimate but faint. Its photometric redshift of places it on the boundary where the Lyman break is just entering the F150W bandpass. Between the different catalogues, this object sometimes used the F150W and F200W or F200W and F277W bands for the 5 detection cuts and its F277W detection was 4.95. So there were instances where if the primary redshift solution was slightly high it was not selected on account of being too faint. Due to its faintness, this source lies within the expected low completeness regime (explored in the next subsection).
In brief, sources that are selected as robust, high-z galaxies in one combination of photometric bands, but not in another, tend to be faint and largely maintain a primary high-z solution in all catalogues. The key difference that impacts their ‘robust’ classification is the relative confidence of the high-z solution compared to the next-best-fitting low-z solution, where the addition of medium bands can either increase or decrease this confidence relative to a broad-band-only catalogue.

### v2

@@ Abstract
New, ultra-deep medium-width photometric coverage with JWST’s NIRCam instrument provides the potential for much improved photo-z reliability at high redshifts. In this study, we conduct a systematic analysis of the JADES Origins Field, which contains 14 broad- and medium-width near-infrared bands, to assess the benefits of medium band photometry on high-z completeness and contamination rates. Using imaging reaching AB mag depth, we test how high-z selections differ when images are artificially degraded or bands are removed. In parallel, the same experiments are conducted on simulated catalogues from the JAGUAR semi-analytic model to examine if observations can be replicated. We find sample completeness is high () and contamination low () when in the regime, even without the use of any medium-width bands. The addition of medium-width bands leads to increases in completeness () but multiple bands are required to improve contamination rates due to the small redshift ranges over which they probe strong emission lines. Incidents of Balmer-Lyman degeneracy increase in the regime and this can be replicated in both simulated catalogues and degraded real data. We measure the faint-end of the UV LF at , finding high number densities that agree with previous JWST observations. Overall, medium bands are effective at increasing completeness and reducing contamination, but investment in achieving comparable depths in the blue (m) as achieved in the red is also found to be key to fully reducing contamination from high-z samples.
@@
In the first few years since the successful commissioning of the James Webb Space Telescope (JWST), the study of the first billion years of the Universe’s history has undergone a revolution. Exquisitely deep, high resolution imaging and spectroscopy has resulted in the photometric selection of ’s of high-z galaxies ( e.g. Harikane et al., 2023; Adams et al., 2024; Hainline et al., 2024; Finkelstein et al., 2024; McLeod et al., 2024; Donnan et al., 2024; D’Silva et al., 2025) and spectroscopic confirmation of ’s (e.g. Bunker et al., 2024; D’Eugenio et al., 2024; Price et al., 2024; Mascia et al., 2024; Heintz et al., 2024). These observations have revealed higher number densities of UV luminous galaxies than first expected from extrapolated HST observations (Oesch et al., 2018; Harikane et al., 2022), pushed the redshift frontier to (Curtis-Lake et al., 2023; Wang et al., 2023; Carniani et al., 2024) and discovered previously unknown classifications of sources (e.g. little red dots Labbé et al., 2023; Matthee et al., 2023; Kokorev et al., 2024).
However, even though JWST provides substantially more information on the Universe than we previously had, and follow up spectroscopy has shown photometric selections have been largely successful, there have been at least two cases where robust looking sources at very high redshifts have turned out to be sample contaminants at lower redshift when examined with NIRSpec (Arrabal Haro et al., 2023; Castellano et al., 2025). Degeneracies between the Lyman or Balmer breaks in a galaxy’s spectra and combinations of dust attenuation with the chance alignment of strong emission lines can result in cases where low-z galaxies (in this case ) enter photometric samples of high-z galaxies (See also discussion in Zavala et al., 2023; Meyer et al., 2024; Gandolfi et al., 2025, for dusty sources). While spectroscopy is the most powerful tool to obtain precise and accurate redshift measurements, JWST has begun to produce deep fields exceeding AB magnitudes of 30 in depth. In this faint regime, follow-up spectroscopy becomes increasingly time consuming, bordering infeasible in fact due to the anticipated costs to obtain required signal-to-noise ratios (Giardino et al., 2022). The use of medium-width photometric bands, while not producing the level of detail as spectroscopy, may provide a route to increase the confidence of high redshift solutions for ultra-faint sources with smaller overall costs. In Cycle 2, the accepted General Observer programme PID 3215 (PI: D. Eisenstein Eisenstein et al., 2023b) produced ultra-deep imaging in multiple medium with bands over a region the GOODS-South field now dubbed the JADES Origins Field (JOF), reaching AB magnitudes of 30.3 in broad- and medium-width bands (0.32as diameter apertures), providing a total of 14 NIRCam bands compared to the more standard 7-8 that many other Cycle 1 and 2 programmes utilised.
@@ 3 Galaxy Selection Procedures
A catalogue of the sources in the Jades Origins Field is produced using the SExtractor software in dual image mode. Here, a weighted stack of the red broad band images (F277W, F356W, F444W) is used at the selection band. Aperture photometry is measured in both 0.2 and 0.32 arcsecond apertures. A set of science images PSF homogenised to the F444W band are also produced using empirical PSF models generated following the technique used in Weaver et al. (2024). We find the profile of the PSF model is similar in shape but slightly broader than theoretical models generated by WebbPSF (Perrin et al., 2012; Perrin et al., 2014). In the following subsections we describe how we select and measure the galaxies using different criteria and later how the results depend on these.
@@ 3.1 Photometry and Photo-Z procedure
Photometric redshifts for our sources are conducted with the use of the EAZY-py SED fitting tool (Brammer et al., 2008). Here, we employ the use of the default FSPS templates as well as Sets 1 and 4 generated in the work by (Larson et al., 2023). These additional templates provide galaxy spectra which are younger, bluer and have stronger emission lines compared to the default template set on its own. EAZY is run using the Madau (1995) treatment for absorption by neutral Hydrogen, adds dust attenuation up to 3.5 magnitudes from Calzetti et al. (2000) and employs no galaxy luminosity prior. EAZY is run twice with upper redshift limits of 25 and 6 in order to conduct a direct comparison between the best fitting high-z and low-z solutions. For our photo-z’s, we quote the peak probability redshift as the primary solution and the 16th/84th percentiles as the error. We also investigate how our results would change when using different size apertures.
@@ 3.3 The initial full sample of high-z candidates
Utilising the full set of NIRCam data available in this field and conducting the above photo-z selection procedure, we obtain 62 galaxies which we class as robust at using the full JOF dataset. The distribution of sources is as follows, with a full table of high-z candidates presented in Table LABEL:tab:fulllist: 30 sources with , 18 sources with , 9 sources with , 4 sources with and 1 source with . Redshifts and values for the sample are shown in Figure 4. Since we only utilise JWST data in our selections, our sample is incomplete at due to the demand for a complete break to be observed and the lack of data bluewards of 0.9 microns. While HST data is present in this field, it is significantly shallower (mag 28.3-29.1 in F606W) and non-uniform (Whitaker et al., 2019). Since ultra-deep HST data can not always be a guarantee for such JWST fields, we focus on the use of just JWST photometry for the purpose of this study. Below we discuss some of the details of our selected galaxies and comparisons to previous similar studies on this field.
@@ 3.3.1 A robust candidate galaxy at
The highest redshift, robust galaxy in our catalogues that passes visual inspection is JADES+53.10762-27.86013, which is a faint () galaxy selected at (See Figure 5) with a of 15 relative to a low redshift solution at . This agrees with the expectations set out in the JOF survey paper (Eisenstein et al., 2023b). Key for determining this is the slightly blue slope measured in the detected bands and the lack of line emission in the F250M band, which covers H and O[III] at . This object is also identified in both JADES photo-z selection papers covering the field, Hainline et al. (2024) and Robertson et al. (2024), as a robust candidate with and respectively. NIRSpec spectroscopy was taken of this source, but due its faint luminosity and the large slit losses in the observations, no definitive redshift was obtained (Carniani et al., 2024).
@@
| JADES+53.09731-27.84714 | Robust | 11.53 | Faint, Broad PDF, Balmer Break (0.2as) | ||
@@
| JADES+53.04017-27.87603 | Robust | 12.10 | Faint (0.32), Robust (0.2) | ||
| JADES+53.03547-27.90037 | Robust | 12.38 | Faint (0.32), Robust (0.2) | ||
| JADES+53.06475-27.89024 | Robust | 12.93 | High (0.32), Robust (0.2) | ||
@@
| JADES+53.08294-27.85563 | Robust | 14.39 | Balmer Break (0.32), Blended, blue (0.2) (Spec-z 14.18) | ||
@@
| JADES+53.07248-27.85535 | Contributors | 10.61 | Robust | ||
| JADES+53.03139-27.87219 | Contributors | 10.76 | Faint | ||
| JADES+53.09292-27.84607 | Contributors | 11.05 | Faint (0.32), 1.6 micron bump (0.32), fail (0.2), Masked | ||
@@
| JADES+53.07385-27.86072 | Secondary | 13.06 | Balmer Break, Faint |
@@ 3.4 Comparisons to the sample selected in Robertson+24
For a full breakdown of our findings for the sources presented in Robertson et al. (2024), we find that: 4/9 sources in their main sample are selected in our work, 0/4 of their lower redshift sources are selected and 0/2 of their dubious sources are selected. In total, four sources are identified in our work to be at and better fit by Balmer breaks than Lyman-breaks. These include both of the sources described as dubious in the work by Robertson et al. (2024). The additional two include one of their lower-z candidates and the spectroscopically confirmed JADES-GS-z14-0 ( Carniani et al., 2024; Schouws et al., 2025; Carniani et al., 2025), which is blended with a neighbour resulting in detections in the bluest bands. We find one source is better fit at as opposed to , however this source has a redshift PDF with a large positive wing and so it is not in significant large tension with the initial Robertson et al. (2024) estimations.
The small area of the JOF field provides some limitations to our work. Firstly, the impact of source blending and artefacts can normally be accounted for when employing the use of large survey areas by removing masked areas/segmentation maps from any area/volume calculations. But for small areas, cosmic variance and small number counts combined with poor luck can result in e.g. one of two-to-three candidates being lost to blending within a masked area of 10%, resulting in underestimations of number density. Whilst the precise aperture location for JADES-GS-z14-0 could be shifted to reduce the flux from the neighbour, the application of this to larger area programmes would become increasingly impractical. Secondly, the small area of the JOF will limit the range of potential SED shapes that are explored. Subsequently, rare or clustered subpopulations that are more or less likely to be mis-modelled in SED fitting could be under- or overrepresented.
@@
Various studies targeting high redshift samples have primarily made use of two different sized circular apertures when conducting initial flux measurements. These have been 0.3-0.32 arcsecond diameters (amounting to 5-6 pixels in radius) or 0.2 arcsecond diameters (amounting to 3.3 pixels in radius). There are potential advantages and disadvantages to using each. The use of larger apertures is slightly more forgiving for the precision of the WCS alignment between photometric bands and it encloses more flux from the PSF (), requiring smaller aperture corrections. The smaller aperture size can enable higher SNR measurements to be made but can be more greatly influenced by bad pixels, correlated noise, minor image misalignments and it requires larger corrections for flux outside of the apertures. At a radius of 0.1 arcsec, the PSF for JWST’s reddest bands has a very steep slope for its encircled flux verses radius, meaning precise PSF models are required as small changes to the FWHM will result in larger changes in the aperture correction verses the use of a larger aperture.
To assess if there are significant differences in sample selection as a consequence of this choice, we repeat our selection procedure on a catalogue using 0.2 arcsecond apertures. This catalogue selects 130 sources at compared to the 62 sources that the 0.32as catalogue did. Of these sources, 52 are consistent between the two samples. for the 10 sources lost, half (5) switch to dominant low-z solutions due to detections in F090W whilst 4 sources have the reduced below constraints and 1 source scatters below the cut.
@@
Among the faint sources added to the high-z sample are an additional four galaxies from the Robertson et al. (2024) sample. Non-selected galaxies from that work with high-z solutions typically only just fail our detection criteria (e.g. 4.5 in rest-frame UV). The spectroscopically confirmed JADES-GS-z14-0 now obtains a high-z solution but remains non-selected on account of contamination from the neighbouring source producing a 3.1 detection in F150W. The source JADES+53.10762-27.86013 (Figure 5) remains the most robust ultra-high-z source in the sample and it is selected with both choices of aperture size. With the smaller aperture size, the to a low-z solution increases from 15 to 22 though the photo-z decreases from 14.7 to 14.0.
@@ 4.1 Adding and Removing Bands from Real Data
Our original selection using the full dataset (C-ALL) identified 62 galaxies at . When applying our selections to the three catalogues with fewer medium bands (C-0, C-1, C-2), each respectively identified 68, 69, 71 candidates. Of these objects, 49, 53, 53 objects are consistent with the fiducial sample (C-ALL), showing that sample overlap with C-ALL is relatively high (79-86%). The largest increase in sample overlap occurs when a single medium-band is added to the broad-band-only dataset. Sources which are identified as confident in C-ALL but omitted when fewer bands are used are largely consistent between the different band combinations too (e.g. of sources selected in C-ALL but omitted in C-2 were also omitted in C-1 and C-0).
We begin our analysis by examining the 13 unique objects that were selected by the C-ALL catalogue yet omitted in the C-0 catalogue. Nine of these sources are low SNR, with F200W detections at . The remaining four are between 8 and 12. Additionally, most of these sources (10) are at the lower end of the redshift range that we consider () where we expect completeness to decrease due to the absence of HST data below m in this study. For the three higher-z galaxies at (all fainter than ), in C-0 we find one galaxy switches to a low-z dominant solution and two fail cuts to the next best low-z solution. In these cases, the medium-width bands in the blue module show a smooth UV slope just before the break, indicating that the break is not a consequence of emission line contamination in the broad-bands.
@@
We next examine sources that are added to the high-z sample when fewer bands are used. For the C-0, C-1 and C-2 catalogues we find that 19, 16, 18 new sources are selected compared to the C-ALL sample. These extra sources are largely consistent between these catalogues, with 13/16 and 14/18 sources in the C-1 and C-2 catalogues appearing in the C-0 catalogue. There are a total of 22 unique sources which get selected as high-z when fewer photometric bands are available. We find that the sources that are selected outside of the C-ALL catalogue are generally faint, with 16/22 sources below in the expected rest-frame UV.
When examining these new sources in the C-ALL catalogue, we find 12/22 sources have primary photo-z solutions that are low-z. These sources can be broken down into 3 categories; 2/12 sources are better fit with 1.6m bumps with all the medium-width photometry available, 2/12 sources are better fit by a Balmer break with no strong emission lines and 8/12 sources are better fit by a Balmer break with strong emission lines. We display an example of each in Figure 6. The strong line emitter category includes three sources exhibiting SED shapes similar to that found in Arrabal Haro et al. (2023) which followed up a candidate (see also Donnan et al., 2023). The use of medium-width bands bluewards of 3 microns is found to be essential to identifying sharp colour differences indicative of emission line presence.
We find that the remaining 10/22 sources lost in C-ALL retain high-z primary solutions but do not meet cuts relative to low-z solutions or fall below the selection. All but one of these galaxies had high-z solution of . Three sources are confident sources that scatter below the cuts for the sample selection. The remainder all contain poorly fit blue medium-width bands (F162M, F182M, F210M) which show colours of up to 1mag and could be indicative of poorly modelled emission lines. The one source at has F182M and F210M photometry showing a smoother break than would otherwise be expected from a Lyman break, reducing its confidence relative to the Balmer break alternative.
In brief, sources that are selected as robust, high-z galaxies in one combination of photometric bands, but not in another, fall within two broad categories. 1) Objects where bluer medium-width bands (<3m) confidently identify or rule-out a low-z solution; 2) Objects where bluer medium-width band observations can not be well described by the template sets used, resulting in poorer quality fits. In cases where confident low-z solutions can be attributed, the most common source of confusion is Balmer break sources with strong emission lines, which represent 2/3rds of potential contaminants.